OpenStudy (anonymous):

whats the answer to: 5-m/3m-15

OpenStudy (anonymous):

-m^2/3-10

OpenStudy (anonymous):

-3?

OpenStudy (anonymous):

-1/3**

OpenStudy (anonymous):

hoqw? how?

OpenStudy (anonymous):

(5-m)/(3*(m-5))= -(m-5)/3*(m-5)= -1/3